# Bayesian Optimization for Non-Convex Two-Stage Stochastic Optimization Problems

### v1

@@ 2.4 Discrete approximation of knowledge gradient
Optimizing the knowledge gradient on line 4 of Algorithm 1 is a non-trivial task Ungredda et al. (2022).
One long-standing method is to replace the search space for the inner optimizations with a finite approximation, ${{\mathbb{X}}_{\mathrm{disc}}}\subset{\mathbb{X}}$. Then the outer expectation can either be computed analytically (Frazier et al., 2009, Algorithm 2) or via a quasi-Monte Carlo approximation, which is the approach taken in this paper.
Write $\mu^{n+1}({\bm{x}}^{\prime};{\bm{x}},v)$ for the posterior mean after $n+1$ samples, where the $(n+1)$th sample is $v$ and made at location ${\bm{x}}$. Conditional on the $n$ observations made so far, the $(n+1)$th observation, ${\textnormal{v}}^{n+1}=f({\bm{x}})+\varepsilon^{n+1}$, is normally distributed, ${\textnormal{v}}^{n+1}\,|\,{\bm{x}}^{1},v^{1},\dots,{\bm{x}}^{n},v^{n}\sim\mathcal{N}(\mu^{n}({\bm{x}}),k^{n}({\bm{x}},{\bm{x}})+\sigma^{2})$. Let $\tilde{v}^{n+1,1},\dots,\tilde{v}^{n+1,N_{v}}$ be a quasi-Monte Carlo sample of ${\textnormal{v}}^{n+1}$, of size $N_{v}$. Then
$\alpha_{\text{KG}}^{n}({\bm{x}})\approx\hat{\alpha}_{\text{KG}}^{n}({\bm{x}})=\frac{1}{N_{v}}\sum_{i=1}^{N_{v}}\max_{{\bm{x}}^{\prime}\in{{\mathbb{X}}_{\mathrm{disc}}}}\mu^{n+1}({\bm{x}}^{\prime};{\bm{x}},\tilde{v}^{n+1,i})-\max_{{\bm{x}}^{\prime}\in{{\mathbb{X}}_{\mathrm{disc}}}}\mu^{n}({\bm{x}}^{\prime}).$ (6)
Using the reparametrization trick Kingma & Welling (2014); Wilson et al. (2017), we can write this as
@@
where ${\textnormal{z}}\sim\mathcal{N}(0,1)$. Fixing $N_{v}$ base samples, $\tilde{z}^{1},\dots,\tilde{z}^{N_{v}}$, of z allows us to differentiate the samples $\{v^{n+1,i}\}_{i}$ in Equation 6 by the input ${\bm{x}}$. Therefore, we can optimize the knowledge gradient with a deterministic, gradient-based optimizer such as multi-start L-BFGS. A Monte Carlo approximation of an expression of the form $\mathbb{E}_{\textnormal{z}}[m({\bm{x}},{\textnormal{z}})]$ is known as sample average approximation (SAA) (Shapiro et al., 2021, Section 5.1).
@@ 3 Knowledge gradient for two-stage problems
We return now to the setting of two-stage stochastic optimization problems, as set out in the problem statement in Section 1.1.
Expected improvement (Jones et al., 1998) is possibly the most common choice of acquisition function in Bayesian optimization. It compares the proposed sample location $({\bm{x}},{\bm{y}},{\bm{u}})$ with the best-tested value so far. However, in two-stage stochastic programming we are searching for a fixed design ${\bm{x}}\in{\mathbb{X}}$ and control policy ${\bm{g}}:{\mathbb{U}}\to{\mathbb{Y}}$, which maximizes the quantity $\bar{f}({\bm{x}},{\bm{g}})=\mathbb{E}_{\mathbf{u}}[f({\bm{x}},{\bm{g}}({\mathbf{u}}),{\mathbf{u}})]$ as in Equation 2. Expected improvement does not readily generalize to this problem because we are making evaluations of $f$ rather than $\bar{f}$ a problem of observations giving partial information. This problem also occurs in robust and multi-task Bayesian optimization, where variants of expected improvement have been proposed (Williams et al., 2000; Groot et al., 2010; Tesch et al., 2011; Ginsbourger et al., 2014). The knowledge gradient acquisition function has also been used to tackle problems of partial information in multi-fidelity and multi-objective optimization and the optimization of function networks (Daulton et al., 2023; Buckingham et al., 2023; Buathong et al., 2024), and is arguably a simpler way to handle partial information. This is the approach taken in this paper.
Given a Gaussian process (GP) model $f\sim\mathcal{GP}(\mu,k)$ of the objective $h$ and its expectation $\overline{f}$ over a random ${\mathbf{u}}$, write $\bar{\mu}^{n}({\bm{x}},{\bm{g}})=\mathbb{E}_{n}[\bar{f}({\bm{x}},{\bm{g}})]$ for the posterior mean of $\bar{f}$ after $n$ observations of $f$. Then, by analogy with Equation 4, after $n$ observations we would recommend a fixed design and control strategy given by
@@
$\alpha_{\text{jKG}}^{n}(\tilde{{\bm{x}}})=\mathbb{E}_{{\textnormal{v}}^{n+1}}\Biggl{[}\underbrace{\max_{\begin{subarray}{c}{\bm{x}}^{\prime}\in{\mathbb{X}}\\ {\bm{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\,\bar{\mu}^{n+1}({\bm{x}}^{\prime},{\bm{g}}^{\prime})}_{\begin{subarray}{c}\text{Maximum of posterior mean}\\ \text{after next observation, }{\textnormal{v}}^{n+1}\end{subarray}}\Bigg{|}\,\tilde{{\bm{x}}}^{n+1}=\tilde{{\bm{x}}}\Biggr{]}\quad-\!\!\!\!\!\underbrace{\max_{\begin{subarray}{c}{\bm{x}}^{\prime}\in{\mathbb{X}}\\ {\bm{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\,\bar{\mu}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})}_{\begin{subarray}{c}\text{Current maximumof posterior}\\ \text{mean (constant wrt }\tilde{{\bm{x}}}\text{)}\end{subarray}}$ (9)
where now ${\textnormal{v}}^{n+1}=f({\bm{x}}^{n+1},{\bm{y}}^{n+1},{\bm{u}}^{n+1})+\varepsilon^{n+1}$ is the next observation (a random variable).
As has been pointed out in previous works, the second term here is constant and can be neglected when optimizing the knowledge gradient. However, as proved in Lemma 3.1, $\alpha_{\text{jKG}}^{n}$ is non-negative, which can be used to improve the choice of initial locations when optimizing it. Therefore, in our experiments, we do not neglect the constant second term.
Lemma 3.1.
@@ 3.1 Asymptotic consistency of the recommendations as estimators of the supremum
It is desirable to know that when using the joint knowledge gradient as an acquisition function, the value, $\overline{h}({\bm{x}}^{*n},{\bm{g}}^{*n})=\mathbb{E}_{\mathbf{u}}[h({\bm{x}}^{*n},{\bm{g}}^{*n}({\mathbf{u}}),{\mathbf{u}})]$, of the solution recommended after $n$ samples converges to the true maximum, $\overline{h}({\bm{x}}^{*},{\bm{g}}^{*})=\mathbb{E}_{\mathbf{u}}[h({\bm{x}}^{*},{\bm{g}}^{*}({\mathbf{u}}),{\mathbf{u}})]$, as the number of samples tends to infinity. Bect et al. (2019) established almost sure convergence with knowledge gradient in the vanilla Bayesian optimization case assuming that the hyperparameters of the Gaussian process surrogate model remain fixed. With the same assumption on the hyperparameters, Toscano-Palmerin & Frazier (2022) showed convergence of the expected values in the risk-neutral, single-stage, stochastic setting if the environmental variable space, ${\mathbb{U}}$, is finite. The following consistency result extends that found in (Bect et al., 2019) to the case of two-stage stochastic optimization.
@@
theoremjointkgconsistency Suppose that the objective, $h$, is a sample from a Gaussian process on ${\mathbb{X}}\times{\mathbb{Y}}\times{\mathbb{U}}$, and that ${\mathbb{X}},{\mathbb{Y}},{\mathbb{U}}$ are compact subsets of $\mathbb{R}^{d_{x}},\mathbb{R}^{d_{y}},\mathbb{R}^{d_{u}}$, respectively. Assume further that the surrogate $f\sim\mathcal{GP}(\mu,k)$ has the same hyperparameters as $h$ so that we may identify $f=h$. Then, viewing the recommendations ${\mathbf{x}}^{*n}$ and ${\mathbf{g}}^{*n}$ as random elements depending on the initial sample, the observation noise, and the function $f$, we have
$\overline{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\to\sup_{\begin{subarray}{c}{\mathbf{x}}\in{\mathbb{X}}\\ {\mathbf{g}}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\overline{f}({\mathbf{x}},{\mathbf{g}})\qquad\text{as}\qquad n\to\infty$
@@
The proof of this theorem can be found in Appendix A and is based on the work by Bect et al. (2019) and Buckingham et al. (2023). It begins by showing that the joint knowledge gradient converges almost surely, uniformly to zero. Further, the posterior mean and covariance of $\overline{f}$ are shown to converge uniformly to continuous limits by observing that the posterior mean and covariance of $f$ are martingales in Banach spaces. These two facts are combined to show that the uniform limit, $\overline{\mu}^{\infty}$, of the posterior means, $\overline{\mu}^{n}$, has sample paths which are almost surely equal to the sample paths of $\overline{f}$ (up to a constant). Whence, convergence of the values of the recommendations can be established.
3.2 Bi-level formulation
Performing optimizations over a function space is a daunting task. By making use of the bi-level formulation introduced in Equation 3, we obtain more tractable expressions. Indeed, by exchanging the order of $\mathbb{E}_{{\textnormal{v}}^{n+1}}$ and $\mathbb{E}_{\mathbf{u}}$, and moving the maximization over ${\bm{g}}$ inside the expectation over ${\mathbf{u}}$ in Equation 8, we obtain
${\bm{x}}^{*n}$ $\in\operatorname*{arg\,max}_{{\bm{x}}\in{\mathbb{X}}}\mathbb{E}_{\mathbf{u}}\Bigl{[}\max_{{\mathbf{y}}\in{\mathbb{Y}}}\mu^{n}({\bm{x}},{\mathbf{y}},{\mathbf{u}})\Bigr{]},$ (10a)
$\forall\,{\bm{u}}\in{\mathbb{U}}\quad{\bm{g}}^{*n}({\bm{u}})$ $\in\operatorname*{arg\,max}_{{\bm{y}}\in{\mathbb{Y}}}\mu^{n}({\bm{x}}^{*n},{\bm{y}},{\bm{u}}),$ (10b)
where $\mu^{n}({\bm{x}},{\bm{y}},{\bm{u}})=\mathbb{E}_{n}[f({\bm{x}},{\bm{y}},{\bm{u}})]$ is the posterior mean of the Gaussian process model.
@@ 3.2 Bi-level formulation
3.3 Efficient computation and optimization
Running a nested optimization is computationally expensive and slow. Instead, we propose a computational strategy based on successive discretization and quasi-Monte Carlo (qMC) approximation similar to that in Section 2.4. Specifically, we take discrete approximations ${{\mathbb{X}}_{\mathrm{disc}}}$ and ${{\mathbb{Y}}_{\mathrm{disc}}}$ in the inner maximizations over ${\mathbb{X}}$ and ${\mathbb{Y}}$, and qMC approximations ${{\mathbb{U}}_{\mathrm{MC}}}$ and ${{\mathbb{V}}_{\mathrm{MC}}}$ for the expectations over ${\mathbf{u}}^{\prime}\in{\mathbb{U}}$ and ${\textnormal{v}}^{n+1}=f(\tilde{{\bm{x}}})+\varepsilon^{n+1}\in\mathbb{R}$. The samples of ${\textnormal{v}}^{n+1}$ are taken conditional on the $n$ observations made so far. In order to be able to differentiate the samples $v\in{{\mathbb{V}}_{\mathrm{MC}}}$ by the proposed next query location $\tilde{{\bm{x}}}$, we use the reparametrization trick as described in Equation 7 to write each $v=v(\tilde{{\bm{x}}},z)$ as a function of $\tilde{{\bm{x}}}$ and a corresponding base sample $z\in{{\mathbb{V}}_{\mathrm{MC}}^{\mathrm{base}}}$, where ${{\mathbb{V}}_{\mathrm{MC}}^{\mathrm{base}}}$ is a Sobol’ qMC approximation of a standard normal variable. Analogously to Equation 6, write $\mu^{n+1}({\bm{x}}^{\prime},{\bm{y}}^{\prime},{\bm{u}}^{\prime};\tilde{{\bm{x}}},v(\tilde{{\bm{x}}},z))$ for the posterior mean of $f({\bm{x}}^{\prime},{\bm{y}}^{\prime},{\bm{u}}^{\prime})$ after $n+1$ samples, where the $(n+1)$th sample is $v(\tilde{{\bm{x}}},z)$ and made at location $\tilde{{\bm{x}}}$. Then, writing $N_{v}=|{{\mathbb{V}}_{\mathrm{MC}}}|=|{{\mathbb{V}}_{\mathrm{MC}}^{\mathrm{base}}}|$ and $N_{u}=|{{\mathbb{U}}_{\mathrm{MC}}}|$ for the sizes of the qMC approximation sets, this gives
$\alpha_{\text{jKG}}^{n}(\tilde{{\bm{x}}})\approx\hat{\alpha}_{\text{jKG}}^{n}(\tilde{{\bm{x}}})=\frac{1}{N_{v}}\sum_{z\in{{\mathbb{V}}_{\mathrm{MC}}^{\mathrm{base}}}}\max_{{\bm{x}}^{\prime}\in{{\mathbb{X}}_{\mathrm{disc}}}}\frac{1}{N_{u}}\sum_{{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}}\max_{{\bm{y}}^{\prime}\in{{\mathbb{Y}}_{\mathrm{disc}}}}\mu^{n+1}({\bm{x}}^{\prime},{\bm{y}}^{\prime},{\bm{u}}^{\prime};\tilde{{\bm{x}}},v(\tilde{{\bm{x}}},z))\\ -\underbrace{\max_{{\bm{x}}^{\prime}\in{{\mathbb{X}}_{\mathrm{disc}}}}\frac{1}{N_{u}}\sum_{{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}}\max_{{\bm{y}}^{\prime}\in{{\mathbb{Y}}_{\mathrm{disc}}}}\mu^{n}({\bm{x}}^{\prime},{\bm{y}}^{\prime},{\bm{u}}^{\prime})}_{\text{constant wrt }\tilde{{\bm{x}}}}.$ (12)
@@ 3.3 Efficient computation and optimization
In order to speed up optimization of the acquisition function, it is important to cache some covariances that are not cached by default in GPyTorch at the time of writing. See Appendix B for details on the caching and Appendix C for a description of the choice of starts for the multi-start L-BFGS-B.
@@
${\bm{x}}^{*n}\approx\hat{{\bm{x}}}^{*n}\in\operatorname*{arg\,max}_{{\bm{x}}\in{\mathbb{X}}}\sum_{{\bm{u}}\in{\mathbb{U}}_{\mathrm{MC},\text{rec}}}\max_{{\bm{y}}\in{\mathbb{Y}}}\mu^{n}({\bm{x}},{\bm{y}},{\bm{u}}).$ (13)
@@
Algorithm 2 shows the full Bayesian optimization loop for the joint knowledge gradient acquisition function. Hyperparameters are fitted at each step using maximum a posteriori (MAP) estimates with the exception of the prior mean, which is fitted only on the first iteration to avoid any bias from collecting samples near the optimum. Further details are given in Section E.3.
@@
3:     Fit hyperparameters of GP prior, $f\sim\mathcal{GP}(\mu,k)$, as described in the text
@@
10:  Optimize ${\bm{x}}^{*n_{\mathrm{tot}}}\in\operatorname*{arg\,max}_{\bm{x}}\frac{1}{N_{u,\text{rec}}}\sum_{{\bm{u}}\in{\mathbb{U}}_{\mathrm{MC},\text{rec}}}\max_{{\bm{y}}}\mu^{n_{\mathrm{tot}}}({\bm{x}},{\bm{y}},{\bm{u}})$
11:  Let ${\bm{g}}^{*n_{\mathrm{tot}}}:{\mathbb{U}}\to{\mathbb{Y}}$ given by ${\bm{g}}^{*n_{\mathrm{tot}}}({\bm{u}})\in\operatorname*{arg\,max}_{{\bm{y}}}\mu^{n_{\mathrm{tot}}}({\bm{x}}^{*n_{\mathrm{tot}}},{\bm{y}},{\bm{u}})$
11:  ${\bm{x}}^{*n_{\mathrm{tot}}},{\bm{g}}^{*n_{\mathrm{tot}}}$

### v2

@@ 2.4 Discrete approximation of knowledge gradient
Optimizing the knowledge gradient on Line 4 of Algorithm 1 is a non-trivial task Ungredda et al. (2022). One long-standing method is to replace the search space for the inner optimizations with a finite approximation, ${{\mathbb{X}}_{\mathrm{disc}}}\subset{\mathbb{X}}$. Then the outer expectation can either be computed analytically (Frazier et al., 2009, Algorithm 2) or via a quasi-Monte Carlo approximation, which is the approach taken in this paper.
Write $\mu^{n+1}({\bm{x}}^{\prime};{\bm{x}},v)$ for the posterior mean after $n+1$ samples, where the $(n+1)$th sample is $v$ and made at location ${\bm{x}}$. Conditional on the $n$ observations made so far, the $(n+1)$th observation, ${\textnormal{v}}^{n+1}=f({\bm{x}})+\varepsilon^{n+1}$, is normally distributed, ${\textnormal{v}}^{n+1}\,|\,({\bm{x}}^{1},v^{1},\dots,{\bm{x}}^{n},v^{n})\sim\mathcal{N}(\mu^{n}({\bm{x}}),k^{n}({\bm{x}},{\bm{x}})+\sigma^{2})$. Let $\tilde{v}^{n+1,1},\dots,\tilde{v}^{n+1,N_{v}}$ be a quasi-Monte Carlo sample of ${\textnormal{v}}^{n+1}$ of size $N_{v}$. Then $\alpha_{\text{KG}}^{n}({\bm{x}})$ is approximated by
Using the reparameterization trick (Kingma & Welling, 2014; Wilson et al., 2017), we can write
@@
where ${\textnormal{z}}\sim\mathcal{N}(0,1)$. Fixing $N_{v}$ base samples, $\tilde{z}^{1},\dots,\tilde{z}^{N_{v}}$, of z allows us to differentiate the samples $\{v^{n+1,i}\}_{i}$ in Equation 6 by the input ${\bm{x}}$. Therefore, we can optimize the knowledge gradient with a deterministic, gradient-based optimizer such as multi-start L-BFGS-B. A Monte Carlo approximation of an expression of the form $\mathbb{E}_{\textnormal{z}}[m({\bm{x}},{\textnormal{z}})]$ is known as sample average approximation (SAA) (Shapiro et al., 2021, Section 5.1).
@@ 3 Knowledge gradient for two-stage problems
We return now to the setting of two-stage stochastic optimization problems, as set out in the problem statement in Section 1.1. We consistently use the following notation. A bar such as in $\bar{\mu}({\bm{x}},{\bm{g}})$ will be used to indicate functions or functionals formed by taking an expectation over the environmental variables ${\mathbf{u}}$, a star such as in ${\bm{x}}^{*n}$ or $\mu^{*n}$ will be used to indicate an optimal quantity (using data available at iteration $n$) and a hat such as in $\hat{\alpha}_{\text{jKG}}$, $\hat{{\bm{x}}}^{*n}$ and $\hat{\bar{\mu}}^{*n}$ will denote an approximation. Additionally, the specific symbol $\tilde{{\bm{x}}}=({\bm{x}},{\bm{y}},{\bm{u}})$ will be used to compactly denote a point in $\tilde{{\mathbb{X}}}={\mathbb{X}}\times{\mathbb{Y}}\times{\mathbb{U}}$. Throughout, we will also remind the reader in words of the meanings of the important quantities.
Expected improvement (Jones et al., 1998) is possibly the most common choice of acquisition function in Bayesian optimization. It compares the proposed sample location $({\bm{x}},{\bm{y}},{\bm{u}})$ with the best-tested value so far. However, in two-stage stochastic programming we are searching for a fixed design ${\bm{x}}\in{\mathbb{X}}$ and control policy ${\bm{g}}:{\mathbb{U}}\to{\mathbb{Y}}$, which maximizes the quantity $\bar{f}({\bm{x}},{\bm{g}})=\mathbb{E}_{\mathbf{u}}[f({\bm{x}},{\bm{g}}({\mathbf{u}}),{\mathbf{u}})]$ as in Equation 2. Expected improvement does not readily generalize to this problem because we are making evaluations of $f$ rather than $\bar{f}$ a problem of observations giving partial information. This problem also occurs in robust and multi-task Bayesian optimization, where variants of expected improvement have been proposed (Williams et al., 2000; Groot et al., 2010; Tesch et al., 2011; Ginsbourger et al., 2014). The knowledge gradient acquisition function has also been used to tackle problems of partial information in multi-fidelity and multi-objective optimization, the optimization of integral expressions, and the optimization of function networks (Daulton et al., 2023; Buckingham et al., 2023; Toscano-Palmerin & Frazier, 2022; Buathong et al., 2024), and is arguably a simpler way to handle partial information. Hence, we advocate the use of knowledge gradient for two-stage stochastic Bayesian optimization.
Given a Gaussian process model $f\sim\mathcal{GP}(\mu,k)$ of the objective $h$ and its expectation $\bar{f}$ over a random ${\mathbf{u}}$, write $\bar{\mu}^{n}({\bm{x}},{\bm{g}})=\mathbb{E}[\bar{f}({\bm{x}},{\bm{g}})\,|\,\mathcal{D}^{n}]$ for the posterior mean of $\bar{f}$ after $n$ observations of $f$. Then, by analogy with Equation 4, after $n$ observations we would recommend a fixed design and control strategy given by
@@
$\alpha_{\text{jKG}}^{n}(\tilde{{\bm{x}}})=\mathbb{E}_{{\textnormal{v}}^{n+1}}\Biggl{[}\underbrace{\max_{\begin{subarray}{c}{\mathbf{x}}^{\prime}\in{\mathbb{X}}\\ {\mathbf{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\,\bar{\mu}^{n+1}({\mathbf{x}}^{\prime},{\mathbf{g}}^{\prime})}_{\begin{subarray}{c}\text{Maximum of posterior mean}\\ \text{after next observation, }{\textnormal{v}}^{n+1}\end{subarray}}\Bigg{|}\,\tilde{{\bm{x}}}^{n+1}=\tilde{{\bm{x}}}\Biggr{]}\quad-\!\!\!\!\!\underbrace{\max_{\begin{subarray}{c}{\bm{x}}^{\prime}\in{\mathbb{X}}\\ {\bm{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\,\bar{\mu}^{n}({\bm{x}}^{\prime},{\bm{g}}^{\prime})}_{\begin{subarray}{c}\text{Current maximumof posterior}\\ \text{mean (constant wrt }\tilde{{\bm{x}}}\text{)}\end{subarray}}$ (9)
where now ${\textnormal{v}}^{n+1}=f({\bm{x}},{\bm{y}},{\bm{u}})+\varepsilon^{n+1}$ is the next observation (a random variable).
As has been pointed out in previous works, the second term here is constant and can be neglected when optimizing the knowledge gradient. However, as proved in Proposition 3.1, $\alpha_{\text{jKG}}^{n}$ is non-negative, which can be used to improve the choice of initial locations when optimizing it in Section 3.2. Therefore, in our experiments, we do not neglect the constant second term.
Proposition 3.1.
@@ 3.1 Asymptotic consistency of the recommendations as estimators of the supremum
It is desirable to know that when using the joint knowledge gradient as an acquisition function, the value, $\bar{h}({\bm{x}}^{*n},{\bm{g}}^{*n})=\mathbb{E}_{\mathbf{u}}[h({\bm{x}}^{*n},{\bm{g}}^{*n}({\mathbf{u}}),{\mathbf{u}})]$, of the solution recommended after $n$ samples converges to the true maximum, $\bar{h}({\bm{x}}^{*},{\bm{g}}^{*})=\mathbb{E}_{\mathbf{u}}[h({\bm{x}}^{*},{\bm{g}}^{*}({\mathbf{u}}),{\mathbf{u}})]$, as the number of samples tends to infinity. Bect et al. (2019) established almost sure convergence with knowledge gradient in the vanilla Bayesian optimization case assuming that the hyperparameters of the Gaussian process surrogate model remain fixed. With the same assumption on the hyperparameters, Toscano-Palmerin & Frazier (2022) showed convergence of the expected values in the risk-neutral, single-stage, stochastic setting if the environmental variable space ${\mathbb{U}}$ is finite. The following consistency result extends that found in (Bect et al., 2019) to the case of two-stage stochastic optimization. The space ${\mathbb{U}}$ need not be finite.
@@
theoremjointkgconsistency Suppose that the objective, $h$, is a sample from a Gaussian process on ${\mathbb{X}}\times{\mathbb{Y}}\times{\mathbb{U}}$ with continuous sample paths, and that ${\mathbb{X}},{\mathbb{Y}},{\mathbb{U}}$ are compact subsets of $\mathbb{R}^{d_{x}},\mathbb{R}^{d_{y}},\mathbb{R}^{d_{u}}$, respectively. Assume further that the surrogate $f\sim\mathcal{GP}(\mu,k)$ has the same hyperparameters as $h$ so that we may identify $f=h$. Then, viewing the recommendations ${\mathbf{x}}^{*n}$ and ${\mathbf{g}}^{*n}$ as random elements depending on the initial sample, the observation noise, and the function $f$, we have
$\bar{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\to\sup_{\begin{subarray}{c}{\mathbf{x}}\in{\mathbb{X}}\\ {\mathbf{g}}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\bar{f}({\mathbf{x}},{\mathbf{g}})\qquad\text{as}\qquad n\to\infty$
@@
The proof of this theorem can be found in Appendix A and is based on the work by Bect et al. (2019) and Buckingham et al. (2023). It begins by showing that the joint knowledge gradient converges almost surely, uniformly to zero. Further, the posterior mean and covariance of $\bar{f}$ are shown to converge uniformly to continuous limits by observing that the posterior mean and covariance of $f$ are martingales in Banach spaces. These two facts are combined to show that the uniform limit, $\bar{\mu}^{\infty}$, of the posterior means, $\bar{\mu}^{n}$, has sample paths that are almost surely equal to the sample paths of $\bar{f}$ (up to a constant). Whence, convergence of the values of the recommendations can be established.
3.2 Efficient computation and optimization
Performing optimizations over a function space is challenging. By making use of the bi-level formulation introduced in Equation 3, we obtain more tractable expressions. Indeed, by exchanging the order of $\mathbb{E}_{{\textnormal{v}}^{n+1}}$ and $\mathbb{E}_{\mathbf{u}}$, and moving the maximization over ${\bm{g}}$ inside the expectation over ${\mathbf{u}}$ in Equation 8, we obtain
${\bm{x}}^{*n}\in\operatorname*{arg\,max}_{{\bm{x}}\in{\mathbb{X}}}\mathbb{E}_{\mathbf{u}}\Bigl{[}\max_{{\mathbf{y}}\in{\mathbb{Y}}}\mu^{n}({\bm{x}},{\mathbf{y}},{\mathbf{u}})\Bigr{]}$ (10a)
[…]
where $\mu^{n}({\bm{x}},{\bm{y}},{\bm{u}})=\mathbb{E}[f({\bm{x}},{\bm{y}},{\bm{u}})\,|\,\mathcal{D}^{n}]$ is the posterior mean of the Gaussian process model.
@@ 3.2 Efficient computation and optimization
Running a nested optimization is computationally expensive and slow. Instead, we propose a computational strategy based on successive discretization and quasi-Monte Carlo (qMC) approximation similar to that in Section 2.4. Specifically, we take discrete approximations ${{\mathbb{X}}_{\mathrm{disc}}}$ and ${{\mathbb{Y}}_{\mathrm{disc}}}$ in the inner maximizations over ${\mathbb{X}}$ and ${\mathbb{Y}}$, and qMC approximations ${{\mathbb{U}}_{\mathrm{MC}}}$ and ${{\mathbb{V}}_{\mathrm{MC}}}$ for the expectations over ${\mathbf{u}}^{\prime}\in{\mathbb{U}}$ and ${\textnormal{v}}^{n+1}=f(\tilde{{\bm{x}}})+\varepsilon^{n+1}\in\mathbb{R}$. The samples of ${\textnormal{v}}^{n+1}$ are taken conditional on the $n$ observations made so far. To differentiate the samples $v\in{{\mathbb{V}}_{\mathrm{MC}}}$ by the proposed next query location $\tilde{{\bm{x}}}$, we use the reparameterization trick as described in Equation 7 to write each $v=v(\tilde{{\bm{x}}},z)$ as a function of $\tilde{{\bm{x}}}$ and a corresponding base sample $z\in{{\mathbb{V}}_{\mathrm{MC}}^{\mathrm{base}}}$, where ${{\mathbb{V}}_{\mathrm{MC}}^{\mathrm{base}}}$ is a Sobol’ qMC approximation of a standard normal variable. Analogously to Equation 6, write $\mu^{n+1}({\bm{x}}^{\prime},{\bm{y}}^{\prime},{\bm{u}}^{\prime};\tilde{{\bm{x}}},v(\tilde{{\bm{x}}},z))$ for the posterior mean of $f({\bm{x}}^{\prime},{\bm{y}}^{\prime},{\bm{u}}^{\prime})$ after $n+1$ samples, where the $(n+1)$th sample is $v(\tilde{{\bm{x}}},z)$ and made at location $\tilde{{\bm{x}}}$. Then, writing $N_{v}=|{{\mathbb{V}}_{\mathrm{MC}}}|=|{{\mathbb{V}}_{\mathrm{MC}}^{\mathrm{base}}}|$ and $N_{u}=|{{\mathbb{U}}_{\mathrm{MC}}}|$ for the sizes of the qMC approximation sets, this gives the approximation
$\hat{\alpha}_{\text{jKG}}^{n}(\tilde{{\bm{x}}})=\frac{1}{N_{v}}\sum_{z\in{{\mathbb{V}}_{\mathrm{MC}}^{\mathrm{base}}}}\max_{{\bm{x}}^{\prime}\in{{\mathbb{X}}_{\mathrm{disc}}}}\frac{1}{N_{u}}\sum_{{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}}\max_{{\bm{y}}^{\prime}\in{{\mathbb{Y}}_{\mathrm{disc}}}}\mu^{n+1}({\bm{x}}^{\prime},{\bm{y}}^{\prime},{\bm{u}}^{\prime};\tilde{{\bm{x}}},v(\tilde{{\bm{x}}},z))\\ -\underbrace{\max_{{\bm{x}}^{\prime}\in{{\mathbb{X}}_{\mathrm{disc}}}}\frac{1}{N_{u}}\sum_{{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}}\max_{{\bm{y}}^{\prime}\in{{\mathbb{Y}}_{\mathrm{disc}}}}\mu^{n}({\bm{x}}^{\prime},{\bm{y}}^{\prime},{\bm{u}}^{\prime})}_{\text{constant wrt }\tilde{{\bm{x}}}}.$ (12)
@@
In order to speed up the optimization of the acquisition function, it is important to cache some covariances that are not cached by default in GPyTorch at the time of writing. See Appendix B for details on the caching and Appendix C for a description of the choice of starts for the multi-start L-BFGS-B.
@@
$\hat{{\bm{x}}}^{*n}\in\operatorname*{arg\,max}_{{\bm{x}}\in{\mathbb{X}}}\sum_{{\bm{u}}\in{\mathbb{U}}_{\mathrm{MC},\text{rec}}}\max_{{\bm{y}}\in{\mathbb{Y}}}\mu^{n}({\bm{x}},{\bm{y}},{\bm{u}}).$ (13)
@@
Algorithm 2 shows the full Bayesian optimization loop for the joint knowledge gradient acquisition function. Hyperparameters are fitted at each step using maximum a posteriori (MAP) estimates. Further details are given in Section E.4.
@@
3:     Fit hyperparameters of GP prior, $f\sim\mathcal{GP}(\mu,k)$, using MAP estimates
@@
10:  Optimize $\hat{{\bm{x}}}^{*n_{\mathrm{tot}}}\in\operatorname*{arg\,max}_{\bm{x}}\frac{1}{N_{u,\text{rec}}}\sum_{{\bm{u}}\in{\mathbb{U}}_{\mathrm{MC},\text{rec}}}\max_{{\bm{y}}}\mu^{n_{\mathrm{tot}}}({\bm{x}},{\bm{y}},{\bm{u}})$
11:  Let $\hat{{\bm{g}}}^{*n_{\mathrm{tot}}}:{\mathbb{U}}\to{\mathbb{Y}}$ given by $\hat{{\bm{g}}}^{*n_{\mathrm{tot}}}({\bm{u}})\in\operatorname*{arg\,max}_{{\bm{y}}}\mu^{n_{\mathrm{tot}}}(\hat{{\bm{x}}}^{*n_{\mathrm{tot}}},{\bm{y}},{\bm{u}})$
11:  $\hat{{\bm{x}}}^{*n_{\mathrm{tot}}},\hat{{\bm{g}}}^{*n_{\mathrm{tot}}}$